A certain mountain has an elevation of 19,381 feet. In 1902, the glacier on this peak covered 8 acres. By 2003 this glacier melted to only 1 acre.
(A)Assume that this glacier melted at a constant rate each year.
(B) Find this yearly rate

Answer by Cromlix(4381)

number of years = 2003 - 1902 = 101 years
Quantity = 8 - 1 = 7 acres
7 acres = 304920ft^2
Annual rate = 304920/101 = 3019ft^2 per year.
Hope this helps.
:-)